1. Calculate the pH of a 1.09×10-2 M solution of the decongestant ephedrine hydrochloride if the pKb of ephedrine (its conjugate base) is 3.86.
1
Expert's answer
2020-04-14T01:43:20-0400

"pK_a+pK_b=14\\pK_a=14-3.86=10.14"

pH of a weak acid solution is "-\log[\sqrt{CK_a}]=-\frac12(\log C+\log K_a)=-\frac12 \log C+\frac12pK_a=-\frac12(-1.96)+5.07=6.05"